It takes Larry 8 hours to paint his house. Larry started to paint the first 3 hours. His friend Patrick joined him and they both finished it in 2 hours. How long will it take Patrick to paint the house if he was to do it alone?
I have tried using fractions and the two equation method. However, what gets me is the 3 hours Larry started to paint by himself. Help!

Found 2 solutions by Theo, richwmiller:
Answer by Theo(13342)

It takes Larry 8 hours to paint his house.
Larry paints for 3 hours by himself.
His friend Patrick joins him and they finish in 2 hours.
How long would it take Patrick to paint the house himself?

Larry would take 8 hours to paint the house by himself.

His rate of painting is 1/8 of the house per hour.

This is based on the formula R * T = U

The number of units is equal to R is equal to 1.

The time is equal to 8 hours.

Plug those numbers into the equation to get R * 8 = 1

Divide both sides of that equation by 8 to get R = 1/8.

Plug 1/8 into that formula to get 1/8 * 8 = 1 so this rate is good.

in 3 hours Larry has painted 3/8 of the house.

That leaves 5/8 of the house left to do.

Larry and Patrick work together to finish the rest of the house in 2 hours.

Since they are working together their rates are additive.

Larry's rate is 1/8 of the house per hour.

Patrick's rate of painting the house per hour is represented by x.

The formula is Number of Units = Rate per Unit * Time

The number of units is equal to 1 if they had to do the whole house.

Since they only have to do 5/8 of the house, then the number of units is 5/8.

Since they are working together, their rates are combined.

The combines rate is (1/8 + x)

the amount of time they took is 2 hours.

The formula becomes:

5/8 = (1/8 + x) * 2

Simplify by removing parentheses to get:

5/8 = 2/8 + 2*x

Subtract 2/8 from both sides of the equation to get:

5/8 - 2/8 = 2*x

Combine like terms to get:

3/8 = 2*x

Divide both sides of the equation by 2 to get:

3/16 = x

Since x is Patrick's rate, this means that Patrick can paint 3/16 of the house in 1 hour.

This means that the formula for Patrick to do the whole house himself would be:

R * T = U where R = 3/16 and U = 1.

This formula becomes (3/16) * T = 1

Divide both sides of this equation by 3/16 to get T = 1/(3/16) which is the same as T = 16/3 which simplifies to T = 5 and 1/3 hours.

It would take Patrick 5 and 1/3 hours to paint the house by himself.

The key to this type of problem is the number of units that are left to be done.


Answer by richwmiller(17219)

Larry worked for 3 hours at the rate of 1/8 per hour.
No need for two equations.
Larry already did 3/8 of the job
Pat and Larry will work together for 2 hours, Larry at 1/8 for each hour and Pat at x per hour for 2 hours
3/8+2/8+2/x=1
5/8=2/x=1
2/x=3/8
16=3x
16/3=x
5 1/3=x
So Pat would do the whole house alone in 5 1/3 hours.
Larry had already done 3/8 with pat he did 2/8. So Larry did 5/8 and Pat did 3/8
of the total job.
Larry worked 5 hours total.
5/8+3/8=1